# MAB-Based Channel Scheduling for Asynchronous Federated Learning in Non-Stationary Environments

### v1

@@ III-A Problem Formulation
The goal of this study is to design an effective online channel scheduling strategy for non-stationary environments with unknown CSI. The aim is to enable clients to identify reliable channels, ensuring successful transmission of parameters to the server in federated learning, thereby accelerating convergence and enhancing the generalization performance of the global model. Based on this, the problem is formulated as follows:
$\mathcal{P}:$ $\min_{\{\boldsymbol{\beta}_{t}\}_{t=1}^{T}}F(\boldsymbol{w}_{T})$ (9a)
s. t. $\sum\nolimits_{n=1}^{N}\beta_{i,k}^{t}=1\quad\forall i\in\mathcal{M},\forall t,$ (9a)
$\sum\nolimits_{i=1}^{M}\beta_{i,k}^{t}\leq 1\quad\forall k\in\mathcal{N},\forall t.$ (9b)
where $\beta_{i,k}$ represents that sub-channel $k$ is allocated to client $i$. $(9a)$ means each client is assigned a channel, and $(9b)$ constraints that each channel is assigned to at most one client.
Problem $\mathcal{P}$ remains challenging to solve because the factors influencing the performance of federated learning are still unclear. The CSI is unknown and the channel state exhibits non-stationary distribution. Consequently, we are unsure how to identify optimal channels for parameter transmission. To address the aforementioned challenges, we begin by performing a convergence analysis of the federated learning framework in non-stationary transmission scenarios, emphasizing the key factors that influence model performance. Based on this analysis, we can then design our channel scheduling strategy.
@@ III-B Convergence Analysis
Assumption 2.(Bounded local gradient) For each client $i\in\{1,2,\dots,M\}$, communication round $t\in\{1,2,\dots,T\}$, and local training epoch $e\in\{1,2,\dots,E\}$, the expected squared norm of the stochastic local gradients is uniformly bounded. i.e., $\mathbb{E}{\left\|{\nabla{F_{i}}(\boldsymbol{w}_{i,t}^{e}|\xi_{i,t}^{e})}\right\|^{2}}\leq{G^{2}}$.
@@
$+\alpha_{3}\frac{1-\Omega^{T}}{1-\Omega}$
@@
Remark 2 : To ensure the model converges, we require $\Omega<1$, which implies $\eta<9L$. As $T\to\infty$, the first term $\Omega^{T}\mathbb{E}[F(\boldsymbol{w}_{0})-F(\boldsymbol{w}^{*})]$ tends to 0. However, the other three terms remain, representing the gap between the global model and the optimal model. To reduce this gap, we can improve model performance by decreasing the second and third terms, which means increasing $|S_{t}|$ and decreasing the average AoI of clients, i.e., $\frac{1}{M}\sum_{i=1}^{M}a_{i}(t)$. Next, we will demonstrate in Lemma 1 that the average AoI of clients is inversely related to $|S_{t}|$, enabling a unified analysis of these two factors.
@@
Therefore, the overall network AoI is $\mathbb{E}[AoI_{t}]=\frac{M^{2}}{|S_{t}|}$.
@@ III-C Problem Transformation
As previously mentioned, the model’s performance and convergence speed are influenced by AoI of the clients. Therefore, selecting good channels for gradient transmission is crucial. However, due to the time-varying nature of the channel’s statistical properties and the unknown CSI, existing methods can not be directly applied to solve this issue. To overcome these challenges, our objective is to leverage the MAB-based framework to design an online channel scheduling strategy that ensures reliable gradient transmission for clients, thus minimizing the cumulative AoI of all clients during $T$ rounds of training. Suppose there exists an oracle strategy that has access to the instantaneous CSI, we reformulate the problem as to design a MAB-based channel scheduling strategy to minimize the AoI regret $R_{\pi}(T)$, which is defined as follows:
$\min R_{\pi}(T)=\sum_{i=1}^{M}\sum_{t=1}^{T}\mathbb{E}[a_{i}^{\pi}(t)-a_{i}^{*}(t)].$ (14)
where $a_{i}^{\pi}(t)$ and $a_{i}^{*}(t)$ denote the AoI of client $i$ at the $t$-th round under the channel scheduling strategy $\pi$ and the oracle strategy, respectively. $R_{\pi}(T)$ represents the total AoI regret of strategy $\pi$ over rounds 1 to $T$. The expectation $\mathbb{E}[\cdot]$ is taken with respect to both the randomization over channel configuration and the policy. Without loss of generality, we assumpt that $a_{i}^{\pi}(0)=a_{i}^{*}(0)=1$ for all clients $i\in\{1,2,\cdots,M\}$ in round $t=0$.
Given the dynamic fluctuations of channel statistics across rounds and the probabilistic success of transmissions, quantifying the number of successful client updates per round presents significant challenges. Our scheduling and allocation policy for non-stationary environments aims to minimize the overall AoI, while also incorporating fairness considerations to ensure effective generalization performance.
@@ IV CHANNEL SCHEDULING IN NON-STATIONARY ENVIRONMENTS
In this section, we will model two types of non-stationary channel environments: extremely non-stationary channels and piecewise stationary channels. For each environment, we design channel scheduling strategies to minimize AoI regret, thereby improving the performance of federated learning models. Additionally, we conduct theoretical analysis of the performance of the proposed algorithms.
@@ IV-A Channel scheduling in extremely non-stationary environment
In the case of extremely non-stationary channels, we model the channel scheduling problem in federated learning as a Multi-Player Multi-Armed Bandit (MP-MAB) problem. Each client is viewed as a player, and each non-stationary channel is viewed as an arm, while the central server acts as the coordinator responsible for selecting and allocating sub-channels. Before introducing the channel scheduling strategy, we first theoretically derive the lower bound of AoI regret that can be achieved in such an environment. We first present the following lemma.
@@
To enable the central server to identify optimal channels, we extend the Exp3.S algorithm [34] and propose the M-Exp3 algorithm tailored for federated learning in extremely non-stationary channels. Specifically, the M-Exp3 algorithm treats the $M$ clients as a group of super players and the combinations of $M$ channels as super arms. In each round, the algorithm probabilistically schedules different channel combinations and adjusts the scheduling probabilities for the subsequent round based on the gradient reception results from the server. The detailed steps of the algorithm are outlined in Algorithm 1.
@@
We now analyze the performance of the M-Exp3 algorithm theoretically, as stated in Theorem 3.
@@
Proof: Let $G_{\max}^{1}(T),\dots,G_{\max}^{M}(T)$ denote the number of successful transmissions (Good state of the channel selected) of each client by the oracle policy with $G_{\max}(T)=\sum_{m=1}^{M}G_{\max}^{m}(T)$, and let $G_{\textup{M-Exp3}}^{1}(T),\dots,G_{\textup{M-Exp3}}^{M}(T)$ denote the corresponding number for M-Exp3 for the $M$ players in $T$ rounds of aggregation with $G_{\textup{M-Exp3}}(T)=\sum_{m=1}^{M}G_{\textup{M-Exp3}}^{m}(T)$. From [34, Corollary 3.2],
@@
$\leq\frac{T^{2}(G_{\max}-G_{\textup{M-Exp3}})}{G_{\max}^{j}(G_{\max}^{j}-2\sqrt{e-1}\sqrt{T|C(M,N)|\log|C(M,N)|})}$ (20)
, where $G_{\max}^{j}G_{\textup{M-Exp3}}^{j}=\min_{i\in[M]}G_{\max}^{i}G_{\textup{M-Exp3}}^{i}$. If $G_{\max}^{j}\geq\frac{T}{|C(N,M)|}$ and $T>16(e-1)|C(N,M)|^{3}\log|C(N,M)|$, then
@@
×$\hskip 85.35826pt\times\sqrt{T|C(N,M)|\log|C(N,M)|}$
IV-B Channel scheduling in piece-wise stationary environment
Piece-wise stationary channel is a weakened version of an extremely non-stationary channel, where the channel mean $\mu_{n},\forall n\in\mathcal{N}$ of each channel remains constant over several consecutive rounds but may experience abrupt changes at certain points. When $\mu$ undergoes a sudden change in a round compared to the previous round, we refer to this round as a breakpoint. Therefore, the channel scheduling problem under piece-wise stationary channels can be modeled as a multi-player piece-wise combinatorial MAB problem. We analyze the minimum achievable AoI regret under such channel conditions:
@@ IV-B Channel scheduling in piece-wise stationary environment
With the piece-wise stationary channel scheduling problem modeled as multi-player combinatorial semi-bandits, for any number of users $M$ , for any number of channels $N\geq M$ and for any aggregation rounds $T$, there exists a distribution over the assignment of channel states such that the AoI regret of any policy is $\Omega(\sqrt{\frac{MT}{N}})$.
@@
To identify the optimal set of channels for gradient transmission, we employ the Combinatorial Upper Confidence Bound Algorithm with Generalized Likelihood Ratio (GLR-CUCB) [26]. The details of the algorithm are presented in Algorithm 2. Next, we conduct a theoretical analysis of the GLR-CUCB algorithm. Since the evolution of AoI in each stationary interval under a piecewise stationary channel is similar to that under a stationary channel, we first derive Lemma 3 and Lemma 4, which pertain to multi-channel scheduling under stationary channels. For analytical convenience, we assume an idealized scenario where each client takes turns using the optimal channel during $T$ rounds of training.
@@
In the stationary channel, let $k_{i}(t)$ be the index of the channel scheduled to client $i$ in round $t$, $k_{i}^{*}(t)$ be the index of the optimal channel scheduled to client $i$ in round $t$ by oracle policy, and $a_{i}(t)$ be AoI of client $i$ in round $t$, such that ${\mu_{\min}}=\mathop{\min}\limits_{k\in[N]}{\mu_{k}}$, $c=\frac{{-1}}{{\log\left({\prod\limits_{k=1}^{M}{(1-{\mu_{k}})}}\right)}}$ and $c^{\prime}=Mc$, we have
$\sum\limits_{t=1}^{T}\mathbb{E}[{a_{i}}(t)]\leq Q+\frac{{c^{\prime}\log T+1}}{{{\mu_{\min}}}}+\frac{{c^{\prime}\log T}}{{{\mu_{\min}}}}\mathbb{E}[{\sum\limits_{t=1}^{T}{{\mathbb{I}_{{k_{i}}(t)\neq k_{i}^{*}(t)}}}}]$ (22)
@@
Proof: When $t>Mc\log T$, let $E_{t}$ represent the event: for $t-Mc\log T+1\leq\tau\leq t$, ${k_{i}}(\tau)=k_{i}^{*}(\tau)$ i.e., ${E_{t}}=\bigcap\nolimits_{\tau=t-Mc\log T+1}^{t}{{k_{i}}(\tau)=k_{i}^{*}(\tau)}$. Let $E_{c}^{t}$ denote the event: $E_{t}^{c}=\bigcup\nolimits_{\tau=t-Mc\log T+1}^{t}{{k_{i}}(\tau)\neq k_{i}^{*}(\tau)}$, then
@@
In federated learning under a piecewise stationary channel environment, channels that were previously in good condition may suddenly degrade, leading to failed gradient update transmissions from clients to the server. To address this issue, we use the Generalized Likelihood Ratio (GLR) as the breakpoint detector and the Combinatorial Upper Confidence Bound (CUCB) as the scheduling algorithm[26]. When no breakpoint is detected, CUCB is used for channel scheduling within each stationary interval. During stationary channel conditions, the UCB value from the previous round is incorporated into the decision for each scheduling round. When a breakpoint is detected, the CUCB algorithm is restarted. The GLR-CUCB algorithm is presented in Algorithm 2 and we have the following lemma:
@@
Lemma 4.
Within the $T$ rounds of aggregation, the piecewise stationary channel can be partitioned into $C_{T}$ intervals, with each interval treated as a stationary channel. The expected total number of occurrences of non-optimal scheduling events under the GLR-CUCB strategy is bounded by ${\cal O}(\sqrt{C_{T}NT\log T})$ when $C_{T}$ is known, and by ${\cal O}(C_{T}\sqrt{NT\log T})$ when $C_{T}$ is unknown.
where $\Delta(t)=\sum\nolimits_{k\in{\textbf{k}^{*}}(t)}{{\mu_{k}}}-\sum\nolimits_{k\in\textbf{k}(t)}{{\mu_{k}}}$ and ${\Delta_{\min}}={\min_{t\in[T]}}\Delta(t)$. $R(T)$ denote the accumulated regret of the GLR-CUCB. With the [26, Corollary 4.3], we arrive at Lemma 4.
Theorem 5.
The AoI regret of all players is,
From Lemma 4, we arrive at Theorem 5.
@@
To adapt MAB-based policies to the AoI Bandit setting, we introduce AoI-Aware (AA) variants that incorporate AoI into decision-making. When AoI is below a predefined threshold, the policy follows the original strategy; otherwise, it prioritizes exploitation based on historical observations to mitigate information staleness.
Let $\sum a_{j}(t)$ denote the total AoI across all clients in round $t$, and $\hat{\mu}_{i}(t)$ represent the estimated transmission success probability of channel $i$ up to round $t$. When AoI is high, the server schedules the $M$ channels with the highest historical success rates to minimize AoI. Otherwise, it focuses on exploring the optimal super-arm via the GLR-CUCB algorithm.

### v2

@@ III-A Problem Formulation
The goal of this study is to design an effective online channel scheduling strategy for non-stationary environments with unknown CSI. This strategy enables clients to identify reliable channels, ensuring successful transmission of parameters to the server in federated learning, thereby accelerating convergence and enhancing the generalization performance of the global model. Based on this, the problem is formulated as follows:
$\mathcal{P}:$ $\min_{\{\boldsymbol{\beta}_{t}\}_{t=1}^{T}}F(\boldsymbol{w}_{T})$ (9)
s.t. $\sum\nolimits_{k=1}^{N}\beta_{i,k}^{t}=1,\quad\forall i\in\mathcal{M},\forall t,$ (9a)
$\sum\nolimits_{i=1}^{M}\beta_{i,k}^{t}\leq 1,\quad\forall k\in\mathcal{N},\forall t,$ (9b)
where $\beta_{i,k}=1$ if sub-channel $k$ is allocated to client $i$, and $\beta_{i,k}=0$ otherwise. $(9a)$ means each client is assigned a channel, and $(9b)$ constraints that each channel is assigned to at most one client.
Problem $\mathcal{P}$ remains challenging to solve because the factors influencing the performance of federated learning are still unclear. The CSI is unknown and the channel state exhibits non-stationary distribution. To address the aforementioned challenges, we begin by performing a convergence analysis of federated learning in non-stationary transmission scenarios, emphasizing the key factors that influence model performance. Based on this analysis, we can then design our channel scheduling strategy.
@@ III-B Convergence Analysis
Assumption 2.(Bounded local gradient) For each client $i\in\{1,2,\dots,M\}$, communication round $t\in\{1,2,\dots,T\}$, and local training epoch $e\in\{1,2,\dots,E\}$, the expected squared norm of the stochastic local gradients is uniformly bounded, i.e., $\mathbb{E}{\left\|{\nabla{F_{i}}(\boldsymbol{w}_{i,t}^{e}|\xi_{i,t}^{e})}\right\|^{2}}\leq{G^{2}}$.
@@
$+\alpha_{3}\frac{1-\Omega^{T}}{1-\Omega},$
@@
Remark 2 : To ensure the model converges, we require $\Omega<1$, which implies $\eta<9L$. As $T\to\infty$, the first term $\Omega^{T}\mathbb{E}[F(\boldsymbol{w}_{0})-F(\boldsymbol{w}^{*})]$ tends to 0. However, the other three terms remain, representing the gap between the global model and the optimal model. To reduce this gap, we can improve model performance by decreasing the second and third terms, which means increasing $|S_{t}|$ and decreasing the average AoI of clients, i.e. $\frac{1}{M}\sum_{i=1}^{M}a_{i}(t)$. Next, we will demonstrate in Lemma 1 that the average AoI of clients is inversely related to $|S_{t}|$, enabling a unified analysis of these two factors.
@@
Therefore, the overall network AoI is $\mathbb{E}[AoI_{t}]=\frac{M^{2}}{|S_{t}|}$. This completes the proof.
@@ III-C Problem Transformation
As previously mentioned, the model’s performance and convergence speed are influenced by AoI of the clients. Therefore, selecting good channels for gradient transmission is crucial. However, due to the time-varying nature of the channel’s statistical properties and the unknown CSI, existing methods can not be directly applied to solve this issue. To overcome these challenges, we leverage the MAB-based framework to design an online channel scheduling strategy that ensures reliable gradient transmission for clients, thus minimizing the cumulative AoI of all clients during $T$ rounds of training. Suppose there exists an oracle strategy that has access to the instantaneous CSI, we reformulate the problem as to design a MAB-based channel scheduling strategy to minimize the AoI regret $R_{\pi}(T)$, which is defined as follows:
$\min R_{\pi}(T)=\sum_{i=1}^{M}\sum_{t=1}^{T}\mathbb{E}[a_{i}^{\pi}(t)-a_{i}^{*}(t)],$ (14)
where $a_{i}^{\pi}(t)$ and $a_{i}^{*}(t)$ denote the AoI of client $i$ at the $t$-th round under the channel scheduling strategy $\pi$ and the oracle strategy, respectively. $R_{\pi}(T)$ represents the total AoI regret of strategy $\pi$ over rounds 1 to $T$. The expectation $\mathbb{E}[\cdot]$ is taken with respect to both the randomization over channel configuration and the policy. Without loss of generality, we assume that $a_{i}^{\pi}(0)=a_{i}^{*}(0)=1$ for all clients $i\in\{1,2,\cdots,M\}$ in round $t=0$.
Given the dynamic fluctuations of channel statistics across rounds and the probabilistic success of transmissions, quantifying the number of successful client updates per round presents significant challenges. Our scheduling policy for non-stationary environments aims to minimize the overall AoI, while also incorporating fairness considerations to ensure effective generalization performance.
@@ IV CHANNEL SCHEDULING IN NON-STATIONARY ENVIRONMENTS
In this section, we model two types of non-stationary channel environments: extremely non-stationary channels and piecewise-stationary channels. For each environment, we design channel scheduling strategies to minimize AoI regret. Additionally, we conduct theoretical analysis of the performance of the proposed algorithms.
@@ IV-A Channel scheduling in extremely non-stationary environment
In the case of extremely non-stationary channels, we model the channel scheduling problem in federated learning as a Multi-Player Multi-Armed Bandit (MP-MAB) problem [35]. Each client is viewed as a player, and each non-stationary channel is viewed as an arm, while the central server acts as the coordinator responsible for selecting and allocating sub-channels. Before introducing the channel scheduling strategy, we first theoretically derive the lower bound of AoI regret that can be achieved in such an environment. We first present the following lemma.
@@
To enable the central server to identify optimal channels, we extend the Exp3.S algorithm [36] and propose the M-Exp3 algorithm tailored for federated learning in extremely non-stationary channels. Specifically, the M-Exp3 algorithm treats the $M$ clients as a group of super players and the combinations of $M$ channels as super arms. In each round, the algorithm probabilistically schedules different channel combinations and adjusts the scheduling probabilities for the subsequent round based on the gradient reception results from the server. The detailed steps of the algorithm are outlined in Algorithm 1. We now analyze the performance of the M-Exp3 algorithm theoretically, as stated in Theorem 3.
@@
Proof: Let $G_{\max}^{1}(T),\dots,G_{\max}^{M}(T)$ denote the number of successful transmissions (Good state of the channel selected) of each client by the oracle policy with $G_{\max}(T)=\sum_{m=1}^{M}G_{\max}^{m}(T)$, and let $G_{\textup{M-Exp3}}^{1}(T),\dots,G_{\textup{M-Exp3}}^{M}(T)$ denote the corresponding number for M-Exp3 for the $M$ players in $T$ rounds of aggregation with $G_{\textup{M-Exp3}}(T)=\sum_{m=1}^{M}G_{\textup{M-Exp3}}^{m}(T)$. From [36, Corollary 3.2],
@@
$\leq\frac{T^{2}(G_{\max}-G_{\textup{M-Exp3}})}{G_{\max}^{j}(G_{\max}^{j}-2\sqrt{e-1}\sqrt{T|C(M,N)|\log|C(M,N)|})},$ (20)
where $G_{\max}^{j}G_{\textup{M-Exp3}}^{j}=\min_{i\in[M]}G_{\max}^{i}G_{\textup{M-Exp3}}^{i}$. If $G_{\max}^{j}\geq\frac{T}{|C(N,M)|}$ and $T>16(e-1)|C(N,M)|^{3}\log|C(N,M)|$, then
@@
×$\hskip 85.35826pt\times\sqrt{T|C(N,M)|\log|C(N,M)|}.$
IV-B Channel scheduling in piecewise-stationary environment
Piecewise-stationary channel is a weakened form of extremely non-stationary channel, where the channel mean $\mu_{i},\forall i\in\mathcal{N}$ of each channel remains constant over several consecutive rounds but may experience abrupt changes at certain rounds. When $\mu$ undergoes a sudden change in a round compared to the previous round, we refer to this round as a breakpoint. Therefore, the channel scheduling problem under piecewise-stationary channels can be modeled as a multi-player piecewise combinatorial multi-arm problem[23]. We analyze the minimum achievable AoI regret under such channel conditions.
@@ IV-B Channel scheduling in piecewise-stationary environment
With the piecewise-stationary channel scheduling problem modeled as multi-player combinatorial semi-bandits, for any number of users $M$ , for any number of channels $N\geq M$ and for any aggregation rounds $T$, there exists a distribution over the assignment of channel states such that the AoI regret of any policy is $\Omega(\sqrt{\frac{MT}{N}})$.
@@
Since the evolution of AoI in each stationary interval under a piecewise-stationary channel is similar to that under a stationary channel, we first derive Lemma 3 and Lemma 4, which pertain to multi-channel scheduling under stationary channels. For analytical convenience, we assume an idealized scenario where each client takes turns using the optimal channel during $T$ rounds of training.
@@
In the stationary channel, let $k_{i}(t)$ be the index of the channel scheduled to client $i$ in round $t$, $k_{i}^{*}(t)$ be the index of the optimal channel scheduled to client $i$ in round $t$ by oracle policy, and $a_{i}(t)$ be AoI of client $i$ in round $t$, such that ${\mu_{\min}}=\mathop{\min}\limits_{k\in[N]}{\mu_{k}}$, $c=\frac{{-1}}{{\log\left({\prod\limits_{k=1}^{M}{(1-{\mu_{k}})}}\right)}}$ and $c^{\prime}=Mc$, we have,
$\sum\limits_{t=1}^{T}\mathbb{E}[{a_{i}}(t)]\leq Q+\frac{{c^{\prime}\log T+1}}{{{\mu_{\min}}}}+\frac{{c^{\prime}\log T}}{{{\mu_{\min}}}}\mathbb{E}[{\sum\limits_{t=1}^{T}{{\mathbb{I}_{{k_{i}}(t)\neq k_{i}^{*}(t)}}}}],$ (22)
@@
Proof: When $t>Mc\log T$, let $E_{t}$ represent the event: for $t-Mc\log T+1\leq\tau\leq t$, ${k_{i}}(\tau)=k_{i}^{*}(\tau)$, i.e. ${E_{t}}=\bigcap\nolimits_{\tau=t-Mc\log T+1}^{t}{{k_{i}}(\tau)=k_{i}^{*}(\tau)}$. Let $E_{c}^{t}$ denote the event: $E_{t}^{c}=\bigcup\nolimits_{\tau=t-Mc\log T+1}^{t}{{k_{i}}(\tau)\neq k_{i}^{*}(\tau)}$, then
@@
This completes the proof.
In federated learning under a piecewise-stationary channel environment, channels that were previously in good condition may suddenly degrade, leading to failed gradient update transmissions from clients to the server. To address this issue, we use the Generalized Likelihood Ratio (GLR) as the breakpoint detector and the Combinatorial Upper Confidence Bound (CUCB) as the scheduling algorithm[24]. When no breakpoint is detected, CUCB is used for channel scheduling within each stationary interval. During stationary channel conditions, the UCB value from the previous round is incorporated into the decision for each scheduling round. When a breakpoint is detected, the CUCB algorithm is restarted. The GLR-CUCB algorithm is presented in Algorithm 2 and we have the following lemma.
Lemma 4.
Within the $T$ rounds of aggregation, the piecewise-stationary channel can be partitioned into $C_{T}$ intervals, with each interval treated as a stationary channel. The expected total number of occurrences of non-optimal scheduling events under the GLR-CUCB strategy is bounded by ${\cal O}(\sqrt{C_{T}NT\log T})$ when $C_{T}$ is known, and by ${\cal O}(C_{T}\sqrt{NT\log T})$ when $C_{T}$ is unknown.
where $\Delta(t)=\sum\nolimits_{k\in{\textbf{k}^{*}}(t)}{{\mu_{k}}}-\sum\nolimits_{k\in\textbf{k}(t)}{{\mu_{k}}}$ and ${\Delta_{\min}}={\min_{t\in[T]}}\Delta(t)$. $R(T)$ denote the accumulated regret of GLR-CUCB. With the [24, Corollary 4.3], we arrive at Lemma 4.
Theorem 5.
The AoI regret of all players is,
From Lemma 4, we arrive at Theorem 5. This completes the proof.
@@
To adapt MAB-based policies to the AoI Bandit setting, we introduce AoI-Aware (AA) variants that incorporate AoI into decision-making. When AoI is below a predefined threshold, the policy follows the original strategy; otherwise, it prioritizes exploitation based on historical observations to mitigate information staleness. Let $\sum a_{j}(t)$ denote the total AoI across all clients in round $t$, and $\hat{\mu}_{i}(t)$ represent the estimated transmission success probability of channel $i$ up to round $t$. When AoI is high, the server schedules the $M$ channels with the highest historical success rates to minimize AoI. Otherwise, it focuses on exploring the optimal super-arm via the GLR-CUCB algorithm.